# Toward a holographic realization of the 2+1-flavor QCD phase structure

### v1

@@ Abstract
While lattice QCD has revealed key features of the QCD phase diagram at zero chemical potential [Borsanyi:2013bia, HotQCD:2014kol], a complete and continuous description of QCD thermodynamics and chiral dynamics across varying quark masses and finite baryon densities remains out of reach. In this context, gauge/gravity duality provides a powerful non-perturbative framework for modeling strongly coupled gauge theories [Witten:1998qj, Gubser:1998bc, Maldacena:1997re]. A long-standing challenge in holographic QCD, however, lies in constructing a unified and dynamically consistent model that faithfully reproduces the full QCD phase structure, including both chiral and deconfinement transitions at zero and finite chemical potential [Gubser:2008yx, DeWolfe:2010he, Jarvinen:2011qe, Chelabi:2015cwn, Li:2016smq, Arefeva:2022avn, Critelli:2017oub, Li:2012ay, Fang:2019lsz, PhysRevD.103.086021, Chen:2019rez, Alho:2012mh, Alfano:2025dch, Li:2014hja, Alho:2013hsa, Fu:2024nmw, Cai:2012xh, Cai:2024eqa, Zhang:2025vsm, Grefa:2022sav, Zhao:2022uxc]. In particular, capturing the quark mass dependence encoded in the Columbia plot remains a key benchmark for the success of any such approach.
@@
which is compared with earlier holographic and QCD-based results [Cai:2022omk, Gao:2020qsj, Li:2018ygx, Zhang:2017icm, Clarke:2024ugt, Shah:2024img, Grefa:2021qvt, Hippert:2023bel]. We also display the experimentally extracted freeze-out line [Lysenko:2024hqp], with the predicted CEP located above it. The experimental coverage shown in Fig. 5 suggests that the predicted CEP lies within reach of upcoming heavy-ion collision programs [Luo:2020pef].
Figure 5: QCD phase diagram from the EMDf model. Dashed and solid lines represent the crossover and first-order transition lines, respectively, with the CEP indicated by a red dot. Prior theoretical estimates of the CEP are included [Cai:2022omk, Gao:2020qsj, Li:2018ygx, Zhang:2017icm, Clarke:2024ugt, Shah:2024img, Grefa:2021qvt, Hippert:2023bel]. The freeze-out line is depicted by gray dotted lines [Lysenko:2024hqp], while top bars highlight the chemical freeze-out regions relevant to heavy-ion collision experiments [Luo:2020pef].

### v2

@@ Abstract
While lattice QCD has revealed key features of the QCD phase diagram at zero chemical potential [Borsanyi:2013bia, HotQCD:2014kol], a complete and continuous description of QCD thermodynamics and chiral dynamics across varying quark masses and finite baryon densities remains out of reach. In this context, gauge/gravity duality provides a powerful non-perturbative framework for modeling strongly coupled gauge theories [Witten:1998qj, Gubser:1998bc, Maldacena:1997re]. A long-standing challenge in holographic QCD, however, lies in constructing a unified and dynamically consistent model that faithfully reproduces the full QCD phase structure, including both chiral and deconfinement transitions at zero and finite chemical potential [Gubser:2008yx, DeWolfe:2010he, Jarvinen:2011qe, Chelabi:2015cwn, Li:2016smq, Arefeva:2022avn, Critelli:2017oub, Li:2012ay, Fang:2019lsz, PhysRevD.103.086021, Chen:2019rez, Alho:2012mh, Alfano:2025dch, Li:2014hja, Alho:2013hsa, Fu:2024nmw, Cai:2012xh, Cai:2024eqa, Zhang:2025vsm, Grefa:2022sav, Zhao:2022uxc, Li:2025lmp, Demircik:2024aig, Ecker:2025vnb]. In particular, capturing the quark mass dependence encoded in the Columbia plot remains a key benchmark for the success of any such approach.
@@
which is compared with earlier holographic and QCD-based results [Cai:2022omk, Gao:2020qsj, Li:2018ygx, Zhang:2017icm, Clarke:2024ugt, Shah:2024img, Grefa:2021qvt, Hippert:2023bel, Jokela:2024xgz]. We also display the experimentally extracted freeze-out line [Lysenko:2024hqp], with the predicted CEP located above it. The experimental coverage shown in Fig. 5 suggests that the predicted CEP lies within reach of upcoming heavy-ion collision programs [Luo:2020pef].
Figure 5: QCD phase diagram from the EMDf model. Dashed and solid lines represent the crossover and first-order transition lines, respectively, with the CEP indicated by a red dot. Prior theoretical estimates of the CEP are included [Cai:2022omk, Gao:2020qsj, Li:2018ygx, Zhang:2017icm, Clarke:2024ugt, Shah:2024img, Grefa:2021qvt, Hippert:2023bel, Jokela:2024xgz]. The freeze-out line is depicted by gray dotted lines [Lysenko:2024hqp], while top bars highlight the chemical freeze-out regions relevant to heavy-ion collision experiments [Luo:2020pef].